### v2

@@ 5.2 Sophisticated dynamic stability
In this section, we provide the natural extension of the solution concept in Hafalir (2008) to the dynamic environment we study. Because Hafalir (2008) names the solution concept sophisticated expectations, we denote ours by sophisticated dynamic stability. Like CVR-dynamic stability, sophisticated dynamic stability is defined recursively. In what follows, we denote by $\mathrm{SDS}_{t}$ the correspondence that maps economies of length $t$ to their set of sophisticated dynamically stable matchings.
Sophisticated dynamic stability builds on two elements: The first is the rational expectations solution discussed at the end of Section 4. Recall that when agent $k$ chooses to remain unmatched in period $t$, it is as if $k$ induces a new economy from period $t$ onward, where everything is as before, except that now $k$ arrives in period $t+1$. Like the rational expectations solution, sophisticated dynamic stability requires that $k$ never rules out matchings $\overline{m}$ such that from period $t$ onward $\overline{m}$ induces a sophisticated dynamically stable matching in the economy induced by $k$ remaining unmatched in period $t$ (see Equation 5.4 below). The second element is a procedure that guarantees the consistency of the conjectures in sophisticated dynamic stability. Recall that 4.1 shows that the conjectures in the rational expectations solution may fail consistency. The procedure below recursively expands the set of conjectures to guarantee their consistency.
We now proceed to define the conjectures for sophisticated dynamic stability, which we denote by $\varphi_{\mathrm{SDS},t}$. Fix an economy $E^{T}$, and a matching through period $t-1$, $m^{t-1}$. Recall that this induces a continuation economy of length $T-(t-1)$, $E^{T}_{t}(m^{t-1})$. Suppose that for all $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we have defined the set of sophisticated dynamically stable matchings for the economy of length $T-(t-1)$ induced by $k$ arriving in period $t+1$, $\mathrm{SDS}_{T-(t-1)}\left(E_{t}^{T}(m^{t-1})_{\setminus k}\right)$.
@@
$F_{0}(E^{T}_{t}(m^{t-1}),k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{s})_{s=t}^{T}\in\mathrm{SDS}_{T-(t-1)}(E_{t}^{T}(m^{t-1})_{\setminus k})\right\}.$ (5.4)
In words, the set $F_{0}(E^{T}_{t}(m^{t-1}),k)$ consists of all sophisticated dynamically stable matchings for the $T-(t-1)$ economy in which $k$ arrives in period $t+1$. We make two observations. First, $F_{0}(E^{T}_{t}(m^{t-1}),k)$ corresponds to the set of conjectures in the “rational expectations solution” using $\mathrm{SDS}_{T-(t-1)}$. Second, in contrast to the notation so far, Equation 5.4 makes the dependence of the conjectures on the continuation economy explicit. The reason is that the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$ depends on the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus k}$, which in turn depends on the conjectures $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1})_{\setminus k},k^{\prime})$ for $k^{\prime}\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\setminus\{k\}$, and so on. Thus, keeping track of the economy from period $t$ onwards is helpful for clarity.
Given $F_{0}(E^{T}_{t}(m^{t-1}),\cdot)$, let $\mathrm{S}(E_{t,F_{0}})$ denote the set of stable matchings in the one-period economy induced by $\{F_{0}(E^{T}_{t}(m^{t-1}),\cdot):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$. Consider now the following set:
@@
$M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1}):\begin{array}[]{l}\text{(i) }\overline{m}_{t}\in\mathrm{S}(E_{t,F_{0}})\\ \text{(ii) }(\overline{m}_{s})_{s=t+1}^{T}\in\mathrm{SDS}_{T-t}(E_{t+1}^{T}(\overline{m}^{t}))\end{array}\right\}.$
$M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ corresponds to the set of matchings such that (i) in period $t$, agents form a stable matching in the one-shot economy induced by their continuation values $F_{0}(E^{T}_{t}(m^{t-1}),\cdot)$, and (ii) the matching from period $t+1$ onward is sophisticated dynamically stable. In other words, $M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ is the set of candidate matchings from Section 4, when the current conjectures are $F_{0}$ and the continuation solution concept is given by $\mathrm{SDS}_{T-t}$.
Recursively, define for $n\geq 1$ and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$,
@@
This process expands the set of conjectures of those agents who find themselves unmatched in the economy with continuation values in $F_{n-1}(E^{T}_{t}(m^{t-1}),\cdot)$. Note that matchings are only added if the elements in $M^{\star}(m^{t-1},F_{n-1},\mathrm{SDS}_{T-1})$ are not part of the set of conjectures $F_{n-1}$.
The recursion in Equation 5.5 is meant to solve the issue discussed in Section 4: If a candidate matching $m^{\star}$ given current conjectures $F_{n-1}$ and the continuation solution concept $\mathrm{SDS}_{T-t}$ is not a solution, then it is blocked by a pair of agents, where at least one of them prefer their worst conjectured matching to matching together. By adding the matchings in $M^{\star}(m^{t-1},F_{n-1},\mathrm{SDS}_{T-1})$, we ensure agents do not dismiss matchings like $m^{\star}$, whose blocks are not credible.
We define $k$’s conjectures as the limit of the set $F_{n}(E^{T}_{t}(m^{t-1}),k)$, that is:
@@
The set in Equation 5.6 is well defined because the sequence $F_{n}$ is increasing (in the set-inclusion sense) and bounded above by $\pazocal{M}_{T}(m^{t-1},k)$.
We can define the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$,
$\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1}),\cdot)$ as the conjectures and $\mathrm{SDS}_{T-t}$ as the continuation solution concept.Footnote 18Footnote 18Footnote 18To be precise, for a given subset $C\subseteq\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we should define the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus C}$ having defined $\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1})_{\setminus C^{\prime}})$ for all subsets $C^{\prime}\subseteq(\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1}))\setminus C$, where $E^{T}_{t}(m^{t-1})_{\setminus C}$ is the length-$T-(t-1)$ economy in which the agents in $C$ arrive in period $t+1$. However, $E^{T}_{t}(m^{t-1}){\setminus C}$ is yet another economy of length $T-(t-1)$, so this omission is hopefully innocuous. Working recursively in this way, we can define the family of conjectures $\Phi_{\mathrm{SDS}}$ for $E^{T}$. The definition of sophisticated dynamic stability is now immediate:
Definition 11.

### v3

@@ 5.2 Sophisticated dynamic stability
In this section, we provide the natural extension of the solution concept in Hafalir (2008) to the dynamic environment we study. Because Hafalir (2008) names the solution concept sophisticated expectations, we denote ours by sophisticated dynamic stability. Like $\mathrm{CVR}$ dynamic stability, sophisticated dynamic stability is defined recursively. In what follows, we denote by $\mathrm{SDS}_{t}$ the correspondence that maps economies of length $t$ to their set of sophisticated dynamically stable matchings.
Agents’ conjectures in sophisticated dynamic stability are defined recursively, building on two elements. First, like the rational expectations solution, conjectures in sophisticated dynamic stability satisfy a form of coherence in the economy induced by an agent’s decision to not be available to match. Recall that when agent $k$ chooses to remain unmatched in period $t$, it is as if $k$ induces a new economy from period $t$ onward, where everything is as before, except that now $k$ arrives in period $t+1$. Similar to the rational expectations solution, sophisticated dynamic stability requires that $k$’s conjectures include all matchings $\overline{m}$ that from period $t$ onward are sophisticated dynamically stable in the economy induced by $k$ remaining unmatched in period $t$ (see Equation 5.4 below). We know from 4.1 that this property alone does not deliver a solution concept with consistent conjectures. The second element is a procedure that guarantees the consistency of the conjectures in sophisticated dynamic stability, by recursively expanding the set of conjectures to guarantee their consistency.
We now define the conjectures for sophisticated dynamic stability, which we denote by $\varphi_{\mathrm{SDS},t}$. Fix an economy $E^{T}$ and a matching through period $t-1$, $m^{t-1}$. Recall $m^{t-1}$ induces a continuation economy of length $T-(t-1)$, $E^{T}_{t}(m^{t-1})$. Suppose that for all $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we have defined the set of sophisticated dynamically stable matchings for the economy of length $T-(t-1)$ induced by $k$ arriving in period $t+1$, $\mathrm{SDS}_{T-(t-1)}\left(E_{t}^{T}(m^{t-1})_{\setminus k}\right)$.
@@
$F_{0}(E^{T}_{t}(m^{t-1}),k)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1},k):(\overline{m}_{t}\setminus k,(\overline{m}_{s})_{s=t+1}^{T})|_{E_{t}^{T}(m^{t-1})_{\setminus k}}\in\mathrm{SDS}_{T-(t-1)}(E_{t}^{T}(m^{t-1})_{\setminus k})\right\},$ (5.4)
where (i) $\overline{m}_{t}\setminus k$ is a period-$t$ matching on $\overline{A}_{t}\cup\overline{B}_{t}\setminus\{k\}$ that coincides with $\overline{m}_{t}$ on that domain, and (ii) we make the dependence of $F_{0}$ on the continuation economy explicit, which is useful because of the recursive nature of the definition.Footnote 18Footnote 18Footnote 18The set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})$ depends on the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus k}$, which in turn depends on the conjectures $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1})_{\setminus k},k^{\prime})$ for $k^{\prime}\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\setminus\{k\}$, and so on.
In words, the set $F_{0}(E^{T}_{t}(m^{t-1}),k)$ consists of all sophisticated dynamically stable matchings for the $T-(t-1)$ economy in which $k$ arrives in period $t+1$. As the construction below makes clear, $F_{0}(E^{T}_{t}(m^{t-1}),k)$ is always a subset of $k$’s conjectures under sophisticated dynamic stability, and hence, sophisticated dynamic stability satisfies a form of coherence similar to that of the rational expectations solution. In fact, if we defined the conjectures of sophisticated dynamic stability using Equation 5.4, sophisticated dynamic stability would coincide with the rational expectations solution.
Given the collection $\{F_{0}(E^{T}_{t}(m^{t-1}),k):k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})\}$, we can now construct the set of candidate matchings given “conjectures” $F_{0}(\cdot)$ and continuation matchings taken from $\mathrm{SDS}_{T-t}$:
@@
$M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)=\left\{\overline{m}\in\pazocal{M}_{T}(m^{t-1}):\begin{array}[]{l}\text{(i) }\overline{m}_{t}|_{E_{t,F_{0}}}\in\mathrm{S}(E_{t,F_{0}})\\ \text{(ii) }(\overline{m}_{s})_{s=t+1}^{T}|_{E_{t+1}^{T}(\overline{m}^{t})}\in\mathrm{SDS}_{T-t}(E_{t+1}^{T}(\overline{m}^{t}))\end{array}\right\}.$
Suppose now a matching $\overline{m}\in M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$ and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ exist such that $\overline{m}_{t}(k)=k$ and $\overline{m}\notin F_{0}(E^{T}_{t}(m^{t-1}),k)$. In other words, $\overline{m}$ fails property CC when using the sets $F_{0}$ as conjectures. Then, we expand the set $F_{0}$ by adding these missing “candidate” matchings, that is,
$F_{1}(E^{T}_{t}(m^{t-1}),k)=F_{0}(E^{T}_{t}(m^{t-1}),k)\cup\left\{\overline{m}\in M^{\star}(m^{t-1},F_{0},\mathrm{SDS}_{T-1}):\overline{m}_{t}(k)=k\right\}.$ (5.5)
Adding these additional matchings to $F_{0}$ is meant to solve the issue discussed in Section 4: Whenever consistency fails and a “candidate” matching, $\overline{m}\in M^{\star}\left(m^{t-1},F_{0},\mathrm{SDS}_{T-t}\right)$, fails to be a solution, it is blocked by a pair of agents, at least one of which prefers their worst conjectured matching in $F_{0}$ to matching together. By adding the matchings in $M^{\star}(m^{t-1},F_{0},\mathrm{SDS}_{T-1})$, we ensure agents do not dismiss matchings such as $\overline{m}$, whose blocks are not credible.
Because $F_{1}$ includes $F_{0}$, the set of candidate matchings given “conjectures” $F_{1}$, $M^{\star}(m^{t-1},F_{1},\mathrm{SDS}_{T-1})$, may include new “candidate” matchings not in $M^{\star}(m^{t-1},F_{0},\mathrm{SDS}_{T-1})$. Furthermore, these new matchings may also be ruled out by blocks that are not credible given “conjectures” $F_{1}$. To address this possibility, we continue to expand the “conjectures” recursively as follows: for $n\geq 1$ and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, define
@@
This process expands the sets $F_{n-1}$ of those agents who find themselves unmatched in the economy with continuation values determined by $F_{n-1}(E^{T}_{t}(m^{t-1}),\cdot)$. Note matchings are only added if the elements in $M^{\star}(m^{t-1},F_{n-1},\mathrm{SDS}_{T-1})$ are not part of the set $F_{n-1}$.
We define $k$’s conjectures as the limit of the set $F_{n}(E^{T}_{t}(m^{t-1}),k)$; that is,
@@
The set in Equation 5.7 is well defined because the sequence $F_{n}$ is increasing (in the set-inclusion sense) and bounded above by $\pazocal{M}_{T}(m^{t-1},k)$. By definition, each element of $M^{\star}(m^{t-1},\varphi_{\mathrm{SDS},t},\mathrm{SDS}_{T-t})$ satisfies property CC when using the sets $\varphi_{\mathrm{SDS},t}$ as conjectures, so the conjectures do not continue to expand once $\varphi_{\mathrm{SDS},t}$ is reached.
We use 4.1 to illustrate the construction of $\varphi_{\mathrm{SDS}}$:
Example 5.1 (continued).
Recall that in the economy of 4.1 the conjectures in the rational expectations solution fail consistency because the candidate matching, denoted by $m^{\star}$ in the example, is not part of agent $a_{3}$’s conjectures. For that reason, we focus in what follows on describing $a_{3}$’s conjectures under sophisticated dynamic stability.
In this example, for each period-$1$ agent, the set $F_{0}(E^{2},k)$ corresponds to their conjectures under rational expectations. The reason is that, for each $k\in A_{1}\cup B_{1}$, the set of sophisticated dynamically stable matchings and the set of rational expectations solutions for $E^{2}_{\setminus k}$ coincide. In particular, in the economy induced by agent $a_{3}$ remaining unmatched in period $1$, $E^{2}_{\setminus a_{3}}$, the following matching is the unique sophisticated dynamically stable matching:
$m_{a_{3}}=\left(\begin{array}[]{lcl}a_{1}&\_\_&b_{2}\\ \hline\cr a_{2}&\_\_&b_{4}\\ a_{3}&\_\_&b_{3}\\ a_{4}&\_\_&b_{1}\end{array}\right),$
and hence, $F_{0}(E^{2},a_{3})=\{m_{a_{3}}\}$.
The same arguments as those in Section 4 imply the set $M^{\star}(\emptyset,F_{0},\mathrm{SDS}_{1})$ is a singleton, consisting of $m^{\star}$. Note agent $a_{3}$ is the only period-$1$ agent who is unmatched under $m^{\star}$ and $m^{\star}\notin F_{0}(E^{2},a_{3})$. Thus, $F_{1}(E^{2},a_{3})=\{m_{a_{3}},m^{\star}\}$, whereas for $k\neq a_{3}$, $F_{1}(E^{2},k)=F_{0}(E^{2},k)$. Because $a_{3}$ is matched with $b_{3}$ under both $m^{\star}$ and $m_{a_{3}}$, the set of candidate matchings given $F_{1}$, $M^{\star}(\emptyset,F_{1},\mathrm{SDS}_{1})$, equals $M^{\star}(\emptyset,F_{0},\mathrm{SDS}_{1})=\{m^{\star}\}$. Because $m^{\star}\in F_{1}(E^{2},a_{3})$, no matching is added to $a_{3}$’s “conjectures” at this step. We then conclude $\varphi_{\mathrm{SDS},1}(E^{2},a_{3})=\{m^{\star},m_{a_{3}}\}$.
Having defined the conjectures under sophisticated dynamic stability for $E^{T}_{t}(m^{t-1})$, we define its set of sophisticated dynamically stable matchings, $\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1}))$, to be the set of matchings for this economy that satisfy 7 with $\varphi_{\mathrm{SDS},t}(E^{T}_{t}(m^{t-1}),\cdot)$ as the conjectures and $\mathrm{SDS}_{T-t}$ as the continuation solution concept.Footnote 19Footnote 19Footnote 19To be precise, for a given subset $C\subseteq\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, we should define the set of sophisticated dynamically stable matchings for $E^{T}_{t}(m^{t-1})_{\setminus C}$, having defined $\mathrm{SDS}_{T-(t-1)}(E^{T}_{t}(m^{t-1})_{\setminus C^{\prime}})$ for all subsets $C^{\prime}\subseteq(\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1}))\setminus C$, where $E^{T}_{t}(m^{t-1})_{\setminus C}$ is the length-$T-(t-1)$ economy in which the agents in $C$ arrive in period $t+1$. However, $E^{T}_{t}(m^{t-1})_{\setminus C}$ is yet another economy of length $T-(t-1)$, so this omission is hopefully innocuous. Working recursively, we can define the family of conjectures $\Phi_{\mathrm{SDS}}$ for $E^{T}$. The definition of sophisticated dynamic stability is now immediate:
Definition 12 (Sophisticated dynamic stability).
@@
When applied to the economy of 5.1, 12 implies $m^{\star}$ is the unique sophisticated dynamically stable matching. That is, in this example, sophisticated dynamic stability coincides with the rational expectations solution, but different conjectures support $m^{\star}$ as a solution in each case.
@@
Comparison with dynamic stability and $\mathrm{CVR}$-dynamic stability
We conclude this section by comparing sophisticated dynamic stability with the solution concepts discussed in Section 5.1.
Our first observation is summarized in 5.2 below: sophisticated dynamic stability is a refinement of dynamic stability:
Proposition 5.2.
For all $T\in\mathbb{N}$, $\mathrm{SDS}_{T}\subseteq\mathrm{DS}_{T}$.
The proof of 5.2 is in Appendix B. We establish that the conjectures in sophisticated dynamic stability are a subset of those under dynamic stability. Therefore, a sophisticated dynamically stable matching is dynamically stable, but as 5.2 below shows, the opposite may not be true. Because, in this example, the set of $\mathrm{CVR}$ dynamically stable matchings coincides with those that are dynamically stable, 5.2 also illustrates that in some economies, sophisticated dynamic stability is a refinement of $\mathrm{CVR}$ dynamic stability.
Example 5.2 ($\mathrm{SDS}\subsetneq\mathrm{DS}$ and $\mathrm{CVR}\neq\mathrm{SDS}$).
Consider the following economy. Arrivals are $A_{1}=\{a_{1},a_{2},a_{3}\}$, $B_{1}=\{b_{1}\}$, $A_{2}=\emptyset$, $B_{2}=\{b_{2},b_{3}\}$. Preferences are given by:
$\begin{array}[]{ccccccc}a_{1}:&(b_{2},0)&(b_{2},1)&(b_{1},0)&(b_{1},1)&(b_{3},0)&(b_{3},1)\\ a_{2}:&(b_{2},0)&(b_{2},1)&(b_{3},0)&(b_{1},0)&(b_{3},1)&(b_{1},1)\\ a_{3}:&(b_{3},0)&(b_{3},1)&(b_{2},0)&(b_{2},1)&(b_{1},0)&(b_{1},1)\end{array}\hskip 14.22636pt\begin{array}[]{ccccccc}b_{1}:&(a_{1},0)&(a_{2},0)&(a_{1},1)&&&\\ b_{2}:&a_{3}&a_{1}&a_{2}&&&\\ b_{3}:&a_{2}&a_{3}&&&&\end{array}$
The following three matchings are $\mathrm{CVR}$ dynamically stable, and hence, dynamically stable:
$m^{L}=\left(\begin{array}[]{lcl}a_{2}&\_\_&b_{1}\\ \hline\cr a_{1}&\_\_&b_{2}\\ a_{3}&\_\_&b_{3}\end{array}\right)\hskip 28.45274ptm^{C}=\left(\begin{array}[]{lcl}a_{1}&\_\_&b_{1}\\ \hline\cr a_{2}&\_\_&b_{3}\\ a_{3}&\_\_&b_{2}\end{array}\right)\hskip 28.45274ptm^{R}=\left(\begin{array}[]{lcl}a_{1}&\_\_&b_{1}\\ \hline\cr a_{2}&\_\_&b_{2}\\ a_{3}&\_\_&b_{3}\end{array}\right).$
The reason is that the following matching is an element of $\varphi_{\mathrm{DS},1}(\emptyset,k)$ and of $\varphi_{\hat{\mathrm{DS}},1}(\emptyset,k)$ for $k\in\{a_{1},a_{2},b_{1}\}$:
$m_{\emptyset}=\left(\begin{array}[]{lcl}&\emptyset&\\ \hline\cr a_{1}&\_\_&b_{1}\\ a_{2}&\_\_&b_{3}\\ a_{3}&\_\_&b_{2}\end{array}\right).$
Because no one matches in $t=1$ under $m_{\emptyset}$, the additional conditions that $\mathrm{CVR}$ dynamic stability imposes on the conjectures do not rule out matching $m_{\emptyset}$ as an element of the conjectures under $\mathrm{CVR}$ dynamic stability.
Instead, only $m^{L}$ is sophisticated dynamically stable. As we show in Appendix B, $m^{\emptyset}$ is not part of the conjectures of $a_{1}$, because $(a_{2},b_{1})$ block $m_{\emptyset}$ in the two-period economy induced by $a_{1}$ waiting to be matched, $E^{2}_{\setminus\{a_{1}\}}=(A_{1}\setminus\{a_{1}\},B_{1},\{a_{1}\},B_{2})$. In fact, the set $\varphi_{\mathrm{SDS},1}(\emptyset,a_{1})$ is a singleton consisting of $m^{L}$, implying that by waiting to be matched, $a_{1}$ can guarantee the payoff of matching with $b_{2}$.
Although 5.2 shows that in some economies, sophisticated dynamic stability refines $\mathrm{CVR}$ dynamic stability, whether this finding holds more generally is an open question.